# Symbiosis: Multi-Adapter Inference and Fine-Tuning

### v1

@@ 3.8 Privacy for Multi-Tenancy
  3. 3.
    Calculation of actual output.
    $y=(Conv1D(x,W)+Conv1D(n,W)+b)-Conv1D(n,W)$
@@
  • $y_{noisy}$: Noisy output and y: Desired output
  • n: Noise and $n_{effect}$: Effect of noise on output
  • W: Parameters and b: Bias
4 Evaluation
Our evaluation test-bed consists of 8 NVIDIA A100 GPUs each with 80GB of memory. The host has 64 AMD EPYC 7763 CPU cores and 512GB of memory. Table 3 list the models used in the following experiments. We compare Symbiosis with the following popular inference and fine-tuning platforms.
  • Baseline: Method provided by Transformers [36] to fine-tune a single adapter. For fine-tuning an adapter using multiple GPUs, we use FSDP as the baseline.
@@ 4 Evaluation
We use LoRA adapters for the following experiments. However, Symbiosis supports other fine-tuning methods such as IA3, Prefix tuning. Table 2 shows the performance of fine-tuning a LoRA adapter with different configurations (Rank, Fine-tuned layers). As compared to the LoRA rank, addition of fine-tuned layers contributes more to increased latency. In the evaluation below, we use the LoRA3 adapter.
@@
Table 2: Fine-tuning iteration latency of LoRA adapters with Llama2-13B. LoRA 1: (8, [q]), LoRA 2: (64, [q]), LoRA 3: (8, [q, k, v, o]), LoRA 4: (64, [q, k, v, o])
@@
Table 3: Models used in the experiments below. This demonstrates the generality of Symbiosis with Llama, BigCodeGPT, Gemma, GPT architectures.
4.1 Memory Consumption
@@
4.1.1 Single Fine-Tuning Job
@@
4.1.2 Multiple Fine-Tuning Jobs
@@
Figure 8: GPU memory consumption of Symbiosis for a file-tuning of a single rank-8 LoRA adapter. Symbiosis-MO represents memory optimized version.
Figure 9: GPU memory consumption for Llama2-13B, batch size=2 and sequence length=512. The base executor memory footprint remains constant with increasing clients.
Figure 10: Single GPU: Iteration latency for Llama3-1B, batch size=2 and sequence length=512. GPU contention hampers baseline latency more than that of Symbiosis.
4.2 Multi-Adapter Fine Tuning
@@
4.2.1 Single GPU Fine Tuning
@@
4.2.2 Multi GPU Fine Tuning
@@
Figure 11: Single GPU: Token throughput for Llama3-1B, batch size=2 and sequence length=512.
Figure 12: Remote Execution: Iteration latency with batch size=2 and sequence length=512. 1 client GPU and 1 base executor GPU.
Figure 13: Remote Execution: Token throughput with batch size=2 and sequence length=512, 1 client GPU and 1 base executor GPU.
@@
For this 2 GPU experiment, we run all clients on a single GPU and the base executor on another GPU. Figures 13 and  13 show per-iteration latency and throughput with increasing number of fine-tuning clients. For Symbiosis, since the clients communicate across the GPUs, our latency and throughput is worse than local configuration. For Llama2-13B, we also observe increasing communication overhead with increasing clients. The fine-tuning performance of Starcoder2-15B is much worse than Llama2-13B from its larger size (60GB). Also, its 32-bit precision requires order of magnitude longer time for common operations, such as matrix multiplication [4], compared to 16-bit precision. The single GPU (baseline) with Starcoder2-15B also requires 3.3s for a fine-tuning iteration with 310 tokens/s throughput (batch size=2 and sequence length=512).
Figure 14: Sharded local: Iteration latency with Llama2-13B, batch Size=2, sequence length=512
Figure 15: Sharded local: Token Throughput with Llama2-13B, batch size=2, sequence length=512.
Figure 16: Sharded Remote: Throughput for Gemma2-27B, batch size=2. The base executor is sharded across 4 GPUs, the clients are hosted on other 4 GPUs.
Figure 17: Heterogeneous GPUs: Throughput for Llama2-13B, batch size=2. C - Client, B - Base executor. Fast represents 350W GPU, Slow represents 100W GPU.
Figure 18: CPU-GPU Inference: KV Cache size is proportional to the Context (Sequence Length). E.g., 128K context length = 64GB KV Cache for Llama2-7B.
Figure 19: Privacy has little impact on inference performance in over-the-network setup.
@@
4.3 Symbiosis with Heterogeneous Resources
@@
4.3.1 Heterogeneous GPU Fine-Tuning
@@
Figure 20: Token generation throughput 8 inference clients for Llama2-7b, batch size=2 and sequence length of 512.
Figure 21: Token generation throughput 6 inference clients and 2 fine-tuning clients for Llama2-7b, batch size=2 and sequence length of 512.
4.3.2 Inference with Heterogeneous Client
@@
4.4 Mixed Fine-tuning and Inference
@@
4.5 Opportunistic Batching
Table 4 shows the benefit of opportunistic batching for inference requests with different characteristics. Specifically, we create a variety by using different batch sizes and different LoRA adapters, LoRA 1 and LoRA 4 (from Table 2). Such a setting is typical for a service provider, who expects requests from a variety of clients. In the no-lockstep approach, each request can progress through the model as quickly as possible without having to batch with other requests. However, with a large number of requests, such approach causes the execution of different layers by different requests to be serialized, which increases response time and the lack of batching also reduces inference throughput. In lockstep policy, the base executor waits for all the requests for every base model layer. Because of the differences in client-side computation, namely, attention and adapter calculation, this policy suffers from longer wait times. In contrast, Symbiosis’s opportunistic batching waits only for a pre-determined duration for batching and then proceeds with the accumulated requests. Importantly, the requests batched at first layer are not required to be batched again for the following layers, which makes different rate of execution for different clients possible. The wait time can be defined by the service provider to capture the worse tolerable penalty in terms of latency.
@@ 4.5 Opportunistic Batching
Table 4: Average response time of 8 inference jobs (token generation) in remote client configuration. The batch sizes (from 2 to 256) and adapter types vary for different clients.
4.6 Privacy for Multi-Tenancy
@@
5 Conclusions
In Symbiosis, we present an inference and fine-tuning platform. Symbiosis provides an abstraction of a layer-as-a-service thus enabling flexible placement, batching and optimizations of individual model layers. We leverage the abstraction to decouple the client specific state, adapters from the base model. This decoupling enables new use cases, namely, (a) base model as-a-service where a shared base model can serve different clients, (b) Inference and fine-tuning over heterogeneous resources, (c) Privacy preserving multi-tenant platform. Moreover, Symbiosis works out-of-the-box for variety of models in transformer library without requiring any changes to the model code.
@@ Acknowledgment
Thanks to Ka-Ho Chow for the initial prototype.

### v2

@@ 3.8. Privacy for Multi-Tenancy
  2. (2)
    Calculation of noisy output by adding $n$ to input $x$ at client before forwarding it to the base executor.
    $y_{noisy}=Conv1D(x+n,W)+b$
    For Conv1D or nn.Linear, the above can be expanded to,
    $y_{noisy}=Conv1D(x,W)+Conv1D(n,W)+b$
  3. (3)
    Calculation of actual output by removing the effect of noise.
    $y=y_{noisy}-n_{effect}$
  • W: Parameters
  • b: Bias
@@
  • y: Desired output
  • n: Noise
  • $y_{noisy}$: Noisy output
  • $n_{effect}$: Effect of noise on output
Since the base executor has access to the noise that is forwarded for calculating the $n_{effect}$, different unique noise values can be used for different layers. This information is only known to the tenant. With only 2 noise values, correctly guessing the noise value used by the client for all layers (e.g., Llama2-7b contains of 100s of nn.Linear layers) is difficult because of extremely large number of possible combinations. Also note that the privacy preservation itself does not address the security of the client environment, where users can rely on existing methods, such as NVIDIA MIG (mig, 2025), to create a secure environment.
4. Evaluation
Our evaluation test-bed consists of 8 NVIDIA A100 GPUs each with 80GB of memory. The host has 64 AMD EPYC 7763 CPU cores and 512GB of memory. Table 3 list the models used in the following experiments. We compare Symbiosis with the following popular inference and fine-tuning platforms. Note that Symbiosis is not an inference or fine-tuning platform by itself, it derives its performance from the optimizations in the underlying Transformers library.
  • Baseline: Method provided by Transformers (Wolf et al., 2020b) to fine-tune a single adapter. For fine-tuning an adapter using multiple GPUs, we use FSDP as the baseline.
@@ 4. Evaluation
  • vLLM: An inference platform, which allows the base model to be shared across multiple PEFT adapter.
@@
We use LoRA adapters for the following experiments. However, Symbiosis supports other fine-tuning methods such as IA3 and prefix tuning. Table 2 shows the performance of fine-tuning a LoRA adapter with different configurations (Rank, Fine-tuned layers). As compared to the LoRA rank, addition of fine-tuned layers contributes more to increased latency. In the evaluation below, we use the LoRA3 adapter.
@@
Table 2. Fine-tuning iteration latency of LoRA adapters with Llama2-13B. LoRA 1: (8, [q]), LoRA 2: (64, [q]), LoRA 3: (8, [q, k, v, o]), LoRA 4: (64, [q, k, v, o])
@@
Table 3. Models used in the experiments below. This demonstrates the generality of Symbiosis with Llama, BigCodeGPT, Gemma, GPT, GPTBigCode architectures.
4.1. Memory Consumption
@@
4.1.1. Single Fine-Tuning Job
@@
4.1.2. Multiple Fine-Tuning Jobs
@@
Figure 9. GPU memory consumption of Symbiosis for a file-tuning of a single rank-8 LoRA adapter. Symbiosis-MO represents memory optimized version.
Figure 10. GPU memory consumption for Llama2-13B, batch size=2 and sequence length=512. The base executor memory footprint remains constant with increasing clients.
Figure 11. Single GPU: Iteration latency for Llama3-1B, batch size=2 and sequence length=512. GPU contention hampers baseline latency more than that of Symbiosis.
4.2. Multi-Adapter Fine Tuning
@@
4.2.1. Single GPU Fine Tuning
@@
4.2.2. Multi GPU Fine Tuning
@@
Figure 12. Single GPU: Token throughput for Llama3-1B, batch size=2 and sequence length=512.
Figure 13. Remote Execution: Iteration latency with batch size=2 and sequence length=512. 1 client GPU and 1 base executor GPU.
Figure 14. Remote Execution: Token throughput with batch size=2 and sequence length=512, 1 client GPU and 1 base executor GPU.
@@
For this 2 GPU experiment, we run all clients on a single GPU and the base executor on another GPU. Figures 14 and  14 show per-iteration latency and throughput with increasing number of fine-tuning clients. For Symbiosis, since the clients communicate across the GPUs, our latency and throughput is worse than local configuration. For Llama2-13B, we also observe increasing communication overhead with increasing clients. The fine-tuning performance of Starcoder2-15B is much worse than Llama2-13B from its larger size (60GB). Also, its 32-bit precision requires order of magnitude longer time for common operations, such as matrix multiplication (Ahmed et al., 2024), compared to 16-bit precision. The single GPU (baseline) with Starcoder2-15B also requires 3.3s for a fine-tuning iteration with 310 tokens/s throughput (batch size=2 and sequence length=512).
Figure 15. Sharded local: Iteration latency with Llama2-13B, batch Size=2, sequence length=512
Figure 16. Sharded local: Token Throughput with Llama2-13B, batch size=2, sequence length=512.
Figure 17. Sharded Remote: Throughput for Gemma2-27B, batch size=2. The base executor is sharded across 4 GPUs, the clients are hosted on other 4 GPUs.
Figure 18. Heterogeneous GPUs: Throughput for Llama2-13B, batch size=2. C - Client, B - Base executor. Fast represents 350W GPU, Slow represents 100W GPU.
Figure 19. CPU-GPU Inference: KV Cache size is proportional to the Context (Sequence Length). E.g., 128K context length = 64GB KV Cache for Llama2-7B.
Figure 20. CPU-GPU Inference with Multiple Requests: Model=Llama2-7b. The client is located on CPU and the base executor is located on GPU.
@@
4.3. Symbiosis with Heterogeneous Resources
@@
4.3.1. Heterogeneous GPU Fine-Tuning
@@
Figure 21. Privacy has little impact on inference performance in over-the-network setup.
Figure 22. Token generation throughput 8 inference clients for Llama2-7b, batch size=2 and sequence length of 512.
Figure 23. Token generation throughput 6 inference clients and 2 fine-tuning clients for Llama2-7b, batch size=2 and sequence length of 512.
4.3.2. Inference with Heterogeneous Client
Single Request:
@@ 4.3.2. Inference with Heterogeneous Client
Multiple Requests:
Figure 20 shows the inference throughput with multiple batched requests, each with 1K sequence length. With 2 GPUs, the client and base executor are placed on different GPUs. However, the 40GB GPU used here for the client cannot accommodate the KV cache for 24 or more requests. In comparison, even though the CPU-side client suffers from higher request latency, it can accommodate 8X as many requests at 7.5 tokens/s throughput.
4.4. Mixed Fine-tuning and Inference
@@
Configuration Latency
Small & Small 0.30
Small & Large 3.74
Large & Large 6.94
Table 4. vLLM prefill response time for batched inference of small and large requests for Llama2-7b model. Small: sequence length=1, Large: sequence length=512.
4.5. Opportunistic Batching
Table 4 shows lockstep execution of two inference requests executed in the same batch in vLLM. Both requests use adapter LoRA 3 from Table 2. It can be observed that when batching large and small requests together, because of the lockstep execution, the response time of the small request also suffers.
In comparison, Table 5 shows the benefit of opportunistic batching for inference requests with different characteristics. Specifically, we create a variety in clients by using different batch sizes (2 to 256) and different LoRA adapters, LoRA 1 and LoRA 4 (from Table 2). Such a setting is typical for a service provider, who expects requests from a variety of clients. In the no-lockstep approach, each request can progress through the model as quickly as possible without having to batch with other requests. However, with a large number of requests, such approach causes the execution of different layers by different requests to be serialized, which increases response time and the lack of batching also reduces inference throughput. In lockstep policy, the base executor waits for all the requests for every base model layer. Because of the differences in client-side computation, namely, attention and adapter calculation, this policy suffers from longer wait times. In contrast, Symbiosis’s opportunistic batching waits only for a pre-determined duration for batching and then proceeds with the accumulated requests. Importantly, the requests batched at first layer are not required to be batched again for the following layers, which makes different rate of execution for different clients possible. The wait time factor can be configured by the service provider to capture the worse tolerable penalty in terms of latency. In this experiment, the request with 256 batch size needs to wait for 50ms at maximum per-iteration.
@@ 4.5. Opportunistic Batching
Table 5. Average response time of 8 inference jobs (token generation) in remote client configuration. The batch sizes (from 2 to 256) and adapter types vary for different clients.
4.6. Privacy for Multi-Tenancy
@@
5. Conclusions
In Symbiosis, we present an inference and fine-tuning platform. Symbiosis provides an abstraction of a base model as-a-service thus enabling flexible placement, batching and optimizations at a granularity of a model layer. We leverage the abstraction to decouple the client specific state, adapters from the base model. This decoupling enables new use cases, namely, (a) base model as-a-service where a shared base model can serve different clients, (b) Inference and fine-tuning over heterogeneous resources, (c) Privacy preserving multi-tenant platform. Moreover, Symbiosis works out-of-the-box for variety of models in transformer library without requiring any changes to the model code.
@@ Acknowledgment
Thanks to Ka-Ho Chow for his help with the initial prototype.